From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue May 06 2003 - 15:21:20 EDT
------------------------------------------------------------ Date: Tue May 6 15:21:20 2003, Shift: 08:00-16:00 Supervisor: JJG, with: Log book#14 pages 96 - 97 Shift Summary: ------------------------------------------------------------ 1. 10 minutes of beam around 9.am 2. collected 185k events (100k trigger 3, 46k trigger 6) between 13.20 and 14.30. 3. RHIC died with QLI and Corrector power supply problem. 4. 12 deg. 1/8 max. field. setting finished. 5. Moved FFS AND BFS to 4 deg. Set magnets to 1/4 field setting for run to explore rates at this angle as input for decision making. 6. No detector problems. 7. Moderate to few achievements ...... ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Tue May 06 2003 - 15:24:16 EDT